# From Question to Exploration: Can Classic Test-Time Adaptation Strategies Be Effectively Applied in Semantic Segmentation?

### v4

@@ 2.2. Classic TTA Strategies
In this paper, our primary objective is to uncover the unique challenges posed by segmentation TTA under classic strategies and provide some inspirational solutions. To achieve that purpose, we delve into several well-established strategies, including normalization updating (Zhao et al., 2023), teacher-student (TS) scheme (Wang et al., 2022a), test-time augmentation (Aug) (Lyzhov et al., 2020), and pseudo labeling (PL) (Zhang et al., 2023b), all of which have demonstrated their effectiveness in classification TTA.
@@ 2.3. Experimental Setups
To ensure consistent evaluations of various TTA approaches, we conduct empirical studies based on several widely used semantic segmentation datasets, including ACDC (Sakaridis et al., 2021), Cityscapes-foggy (CS-fog) (Sakaridis et al., 2018) and Cityscapes-rainy (CS-rain) (Hu et al., 2019). In addition, we strictly follow the implementation details outlined in previous studies (Wang et al., 2022a; Colomer et al., 2023), and use Segformer-B5 (Xie et al., 2021) as the pre-trained model. Two state-of-the-art and recent segmentation approaches, i.e., Oneformer (Jain et al., 2023) and SAM (Kirillov et al., 2023), are also used in comparative experiments. We focus on transformer-based architectures instead of CNN-based architectures, since the former has exhibited more promising results than the latter. Unless otherwise specified, all experiments are conducted with a batch size (BS) of 1, mirroring real-world scenarios where the test samples often arrive one by one in an online manner. The choice of hyper-parameters can be seen in the code of this paper which is publicly available.
@@ 3.3. Distribution Estimation Tricks Cannot Resolve the Problem
Utilizing local test mini-batch statistics for model adaptation proves unreliable, especially when confronting persistent distribution shifts (Yuan et al., 2023; Zhang et al., 2020; Qiu et al., 2021). Such unreliability originates from error gradients and imprecise estimations of test data statistics. In response, we delve into two test-time batch renormalization techniques (Zhao et al., 2023; Yuan et al., 2023), namely Test Local Adapt and Test Global Adapt, aiming to refine the distribution estimation. Test Local Adapt leverages the source statistics to recalibrate the mini-batch test data distribution estimation, whereas Test Global Adapt uses test-time moving averages to recalibrate the overall test distribution estimation. As shown in Figure 1 (c), while batch renormalization strategies do enhance the performance of TENT, their performance is just comparable to that of SO and cannot lead to performance improvement in semantic segmentation.
@@ 4.2. Discussions of Potential Limitations
While previous analysis attests to the efficacy of the TS-EMA scheme, we find that when the SO baseline is fortified with test-time augmentation, its performance surpasses that of TS combined with both PL and Aug. This suggests that the primary advantage of TS-EMA may lie in mitigating the noise introduced by PL, thereby allowing Aug to function more effectively.
@@
We continue to compare the single-model and the TS scheme. As depicted in Figure 2, we have plotted the IoU (Intersection over Union) metrics for each class against varying levels of GT. This visualization helps us critically assess how the performance trajectory of these two schemes adjusts as the accuracy of the PLs is promoted. For the sake of fair comparison, the policy of Aug is not adopted and it is clear that the performance improvement will be minimal without data augmentation. Our purpose is to investigate the importance of each module of the TS scheme and emphasize the necessity of Aug in this scheme.
Figure 2. Comparisons between the single-model and the teacher-student (TS) scheme under different degrees of ground-truth (GT) PLs (%) on ACDC. As the accuracy of PLs increases, the performance of the single-model experiences continual enhancement. However, the TS scheme’s performance remains stagnant since the strategy of test-time augmentation has not been introduced.
@@ 5.1. Segmentation TTA Suffers Long-tailed Problem
When placed in the context of TTA, the long-tailed (LT) problem becomes more pronounced, manifesting as an obvious bias in test-time optimization towards dominant classes (Zhao et al., 2023; Zhang et al., 2022). Both NOTE (Gong et al., 2022) and SAR (Niu et al., 2023) can handle the class imbalance in classification TTA, however, they perform poorly when addressing the LT problem in segmentation TTA. In this Section, we aim to shed light on the challenges of the LT problem as it manifests in segmentation TTA. We show the intricate complexity and challenge inherent in semantic segmentation, making it markedly more difficult than classification tasks. To delve deeper into this issue, we assume that the model can generate high-confidence PLs for the test data during adaptation and subsequently analyze the resultant state of the model. Our analysis will be conducted from two perspectives: conducting recall-precision analysis, and evaluating model calibration.
@@
To further detailed analysis of LT, we also show the quantitative metrics of each class on ACDC-fog, as shown in Figure 5. We conduct a comparison of the results between two experiments: Source Only (SO) and Adapt (where we fine-tune the source model using 100% GT labels). Firstly, as evident in all the plots of this figure, the majority classes consistently achieve exceptionally high scores across all metrics, whereas the minority classes do not consistently perform the worst. Secondly, following the adaptation process (involving the addition of supervised information to model training), the recall of most classes shows improvement, while the precision of certain minority classes experiences a decrease. This indicates that the model is less likely to miss pixels of this class (predicting it as other classes) while becoming more prone to predicting pixels of other classes as this class. This phenomenon diverges from the patterns observed in classification tasks (Wei et al., 2021) and does not align with conventional wisdom, adding complexity to the uncovering of underlying patterns.
@@ 5.2. How to Relieve LT Biases?
Furthermore, we conduct an ablation study for test-time augmentation (Lyzhov et al., 2020; Wang et al., 2022a; Zhang et al., 2023c) in terms of the F1 Score and mIoU. As shown in Table 5, it is clear that employing data augmentation results in a 2.4% increase in mIoU. However, it simultaneously leads to a 0.9% decrease in the F1 Score. This suggests that the model, post-augmentation, intensifies its prediction of minority classes, leading to a simultaneous rise in both True Positive and False Positive, thereby boosting mIoU. Nonetheless, the nuanced balance of Recall and Precision in the F1 Score leads to a less pronounced change. Regarding the tail classes, we observe a notable 4.4% increase in mIoU, contrasted by a 1.1% decline in F1 Score. This showcases that while augmentation enhances the model’s detection of tail classes, it does not uniformly improve its precision for these classes. In light of the above observations, we conclude that Aug partially relieves LT biases in segmentation TTA. In the future, we will explore integrating region-level segmentation and Aug to address the LT problem in segmentation TTA.
Table 5. Ablation studies on ACDC-fog of data augmentation (Aug) in terms of F1 Score and mIoU (%).
Method Aug F1 Score mIoU
head mid tail Avg. head mid tail Avg.
Pseudo labeling 89.8 82.4 82.7 85.6 82.8 71.1 69.9 74.5
89.7 (-0.1) 82.7 (+0.3) 81.6 (-1.1) 84.7 (-0.9) 82.9 (+0.1) 73.5(+2.4) 74.3(+4.4) 76.9(+2.4)
@@ 7. Conclusions
Besides the regular observations, we discover that visual prompt tuning (VisPT) is a promising solution to address segmentation TTA. Consequently, we propose a novel method named TTAP which has also been proved to be effective. We hope that more researchers can join the TTA community and build a common practice for segmentation.

### v5

@@ 2.2. Classic TTA Strategies
In this paper, our primary objective is to uncover the unique challenges posed by segmentation TTA under classic strategies and provide some inspirational solutions. To achieve that purpose, we delve into several well-established strategies, including normalization updating (Zhao et al., 2023a), teacher-student (TS) scheme (Wang et al., 2022a), test-time augmentation (Aug) (Lyzhov et al., 2020), and pseudo labeling (PL) (Zhang et al., 2023b), all of which have demonstrated their effectiveness in classification TTA.
@@ 2.3. Experimental Setups
To ensure consistent evaluations of various TTA approaches, we conduct empirical studies based on several widely used semantic segmentation datasets, including ACDC (Sakaridis et al., 2021), Cityscapes-foggy (CS-fog) (Sakaridis et al., 2018) and Cityscapes-rainy (CS-rain) (Hu et al., 2019). In addition, we strictly follow the implementation details outlined in previous studies (Wang et al., 2022a; Colomer et al., 2023), and use Segformer-B5 (Xie et al., 2021) as the pre-trained model. Two state-of-the-art and recent segmentation approaches, i.e., Oneformer (Jain et al., 2023) and SAM (Kirillov et al., 2023), are also used in comparative experiments. We focus on transformer-based architectures instead of CNN-based architectures, since the former exhibits more promising results than the latter (cf. Appendix 1). Unless otherwise specified, all experiments are conducted with a batch size (BS) of 1, mirroring real-world scenarios where the test samples often arrive one by one in an online manner. Some of the experimental results, i.e., Tables and Figures, are displayed in the Appendix. The choice of hyper-parameters can be seen in the code of this paper which will be publicly available.
@@ 3.3. Distribution Estimation Tricks Cannot Resolve the Problem
Utilizing local test mini-batch statistics for model adaptation proves unreliable, especially when confronting persistent distribution shifts (Yuan et al., 2023; Zhang et al., 2020; Qiu et al., 2021). Such unreliability originates from error gradients and imprecise estimations of test data statistics. In response, we delve into two test-time batch renormalization techniques (Zhao et al., 2023a; Yuan et al., 2023), namely Test Local Adapt and Test Global Adapt, aiming to refine the distribution estimation. Test Local Adapt leverages the source statistics to recalibrate the mini-batch test data distribution estimation, whereas Test Global Adapt uses test-time moving averages to recalibrate the overall test distribution estimation. As shown in Figure 1 (c), while batch renormalization strategies do enhance the performance of TENT, their performance is just comparable to that of SO and cannot lead to performance improvement in semantic segmentation.
@@ 4.2. Discussions of Potential Limitations
While previous analysis attests to the efficacy of the TS-EMA scheme, a closer examination of Table 6 (cf. Appendix) underscores a notable observation: when the SO baseline is fortified with test-time augmentation, its performance surpasses that of TS combined with both PL and Aug. This suggests that the primary advantage of TS-EMA may lie in mitigating the noise introduced by PL, thereby allowing Aug to function more effectively.
@@
We continue to compare the single-model and the TS scheme. As depicted in Figure 5 (cf. Appendix), we have plotted the IoU (Intersection over Union) metrics for each class against varying levels of GT. This visualization helps us critically assess how the performance trajectory of these two schemes adjusts as the accuracy of the PLs is promoted. For the sake of fair comparison, the policy of Aug is not adopted in that Figure, where comparative results indicate that the performance improvement will be minimal without data augmentation. This experiment aims to investigate the importance of each module of the TS scheme and emphasize the necessity of Aug in this scheme. Moreover, we also report the result of the TS scheme leveraging data augmentation in Figure 6 (cf. Appendix).
@@ 5.1. Segmentation TTA Suffers Long-tailed Problem
When placed in the context of TTA, the long-tailed (LT) problem becomes more pronounced, manifesting as an obvious bias in test-time optimization towards dominant classes (Zhao et al., 2023a; Zhang et al., 2022). Both NOTE (Gong et al., 2022) and SAR (Niu et al., 2023) can handle the class imbalance in classification TTA, however, they perform poorly when addressing the LT problem in segmentation TTA. As shown in Figure 12 (cf. Appendix), the numerical disparity between the majority and minority classes surpasses a staggering 1000-fold difference. This stark contrast is evident when compared to common datasets used in classification tasks, such as CIFAR10-LT, where the most majority class is only in the thousand-level range and has 100$\times$ more samples than the most minority class (Wei et al., 2021). Adding to the challenge is the nature of semantic segmentation itself, which involves copious pixel-level labels, further complicating the LT complexity. In this Section, we aim to shed light on the challenges of the LT problem as it manifests in segmentation TTA.
We then show the intricate complexity and challenge inherent in semantic segmentation, making it markedly more difficult than classification tasks. To delve deeper into this issue, we assume that the model can generate high-confidence PLs for the test data during adaptation and subsequently analyze the resultant state of the model. Our analysis will be conducted from three perspectives: examining the confusion matrix, conducting recall-precision analysis, and evaluating model calibration.
Confusion matrix.
The confusion matrix of ACDC-fog is displayed in Figure 15 (cf. Appendix), unveiling extreme variations in the outcomes for each class, reflecting the substantial discrepancy in the metric across different classes. For example, when a pixel is predicted to be fence, the possibilities of its true labels—rider, motorcycle, and bicycle—are all less than $10^{-6}$, contrasting sharply with other classes that are in the tens of thousands. We suggest this stark difference elucidates the extreme variation and irregularity in the model’s predictive accuracy for different classes.
@@
To further detailed analysis of LT, we also show the quantitative metrics of each class on ACDC-fog [^3] The results on the other domains of dataset ACDC are presented in Figure 9-Figure 11 (cf. Appendix)., as shown in Figure 8 (cf. Appendix). We conduct a comparison of the results between two experiments: Source Only (SO) and Adapt (where we fine-tune the source model using 100% GT labels). Firstly, as evident in all the plots of this figure, the majority classes consistently achieve exceptionally high scores across all metrics, whereas the minority classes do not consistently perform the worst. Secondly, following the adaptation process (involving the addition of supervised information to model training), the recall of most classes shows improvement, while the precision of certain minority classes experiences a decrease. This indicates that the model is less likely to miss pixels of this class (predicting it as other classes) while becoming more prone to predicting pixels of other classes as this class. This phenomenon diverges from the patterns observed in classification tasks (Wei et al., 2021) and does not align with conventional wisdom, adding complexity to the uncovering of underlying patterns.
@@ 5.2. How to Relieve LT Biases?
Furthermore, we consider the individual role of augmentation, and the results are displayed in Table 6 (cf. Appendix), pondering the potential of test-time augmentation to alleviate the issue of tail-class information scarcity (Zhang et al., 2023c). Following this, we conduct an ablation study for test-time augmentation (Lyzhov et al., 2020; Wang et al., 2022a) in terms of the F1 Score and mIoU. As shown in Table 7 (cf. Appendix), it is clear that employing data augmentation results in a 2.4% increase in mIoU. However, it simultaneously leads to a 0.9% decrease in the F1 Score. This suggests that the model, post-augmentation, intensifies its prediction of minority classes, leading to a simultaneous rise in both True Positive and False Positive, thereby boosting mIoU. Nonetheless, the nuanced balance of Recall and Precision in the F1 Score leads to a less pronounced change. Regarding the tail classes, we observe a notable 4.4% increase in mIoU, contrasted by a 1.1% decline in F1 Score. This showcases that while augmentation enhances the model’s detection of tail classes, it does not uniformly improve its precision for these classes. In light of the above observations, we conclude that Aug partially relieves LT biases in segmentation TTA. In the future, we will explore integrating region-level segmentation and Aug to address the LT problem in segmentation TTA.
@@ 7. Conclusions
Besides the regular observations, we discover that visual prompt tuning (VisPT) is a promising solution to address segmentation TTA. Consequently, we propose a novel method named TTAP which has also been proved to be effective. More information such as Tables, Figures, and analysis can be found in the Supplementary Material. We hope that more researchers can join the TTA community and build a common practice for segmentation.
@@ Acknowledgements.
Appendix A Related Studies
Classic test-time adaptation
Normalization statistics are widely used in TTA to compute the data distribution based on the test data. TENT (Wang et al., 2021) adapts batch normalization (BN) layers based on entropy minimization, i.e., the confidence of the target model is measured by the entropy of its predictions. EATA (Niu et al., 2022) actively selects reliable samples to minimize entropy loss during inference. Furthermore, it also introduces a Fisher regularizer to filter out redundant samples to reduce the computational time. SAR (Niu et al., 2023) is a reliable and sharpness-aware entropy minimization approach that can suppress the effect of noisy test samples with large gradients. ATP (Bao et al., 2024) is flexible to handle various kinds of distribution shifts in online federated learning, by adaptively learning the adaptation rates for each target model. However, the cross-entropy loss, which is effectively used in classification, is inherently inapplicable to a regression problem such as pose estimation (Li et al., 2021).
Table 8. Results based on DeepLabv3+, a CNN-based architecture. Compared to the counterpart that is based on Transformer (Table 1 of the main manuscript), the performance drops about 10% in average.
Method A-fog A-night A-rain A-snow CS-fog CS-rain Avg.
SO 58.9 32.0 48.1 45.7 65.9 49.8 50.1
BN adapt 36.2 28.9 37.7 36.1 60.5 55.4 47.5
TENT (Wang et al., 2021) 59.2 31.9 48.8 46.6 64.9 53.1 50.7
Besides entropy-based approaches, many other strategies are also introduced to address TTA. TEA (Yuan et al., 2024) transforms the source model into an energy-based classifier to align the distributions of the model and test data. AdaContrast (Chen et al., 2022) combines contrastive learning and pseudo labeling to handle TTA. AdaNPC (Zhang et al., 2023b) is a parameter-free TTA approach based on a K-Nearest Neighbor (KNN) classifier, where the voting mechanism is used to attach labels based on $k$ nearest samples from the memory. Different from traditional approaches, CTTA-VDP (Gan et al., 2023) introduces a homeostasis-based prompt adaptation strategy that freezes the source model parameters during the continual TTA process. Based on large-scale open-sourced benchmark approaches and thorough analysis, TTAB (Zhao et al., 2023b) unveils three pitfalls in prior TTA approaches under classification tasks.
Semantic segmentation
Pixel-level annotation is one of the key characteristics of semantic segmentation. HAMLET (Colomer et al., 2023) can handle unforeseen continuous domain changes since it combines a specialized domain-shift detector and a hardware-aware backpropagation orchestrator to actively control the model’s real-time adaptation for semantic segmentation. CoTTA (Wang et al., 2022a) can reduce error accumulation based on weight-averaged and augmentation-averaged predictions. Segmentation tasks are also pervasive in medical images since the scanner model and the protocol differ across different hospitals. This issue can be handled by introducing an adaptable per-image normalization module and denoising autoencoders to incentivize plausible segmentation predictions (Karani et al., 2021).
SITA (Khurana et al., 2023) can be applied in segmentation and the source model is adapted independently based on each individual test sample which will be augmented several times. DIGA (Wang et al., 2023b) is a backward-free segmentation approach that is based on a semantic and a distribution adaptation module, which can adapt the model at both semantic and distribution levels. However, the weights of different modules are fixed. Segmentation TTA has also been extended to multi-modal 3D tasks based on intra-modal pseudo-label generation and inter-modal pseudo-label refinement (Shin et al., 2022), although the experiments are carried out on simple scenarios. OASIS (Volpi et al., 2022) is a training-validation-deploy benchmark that focuses on the evaluation protocol, adaptation benchmark, and impact of catastrophic forgetting.
Similar to TTAB (Zhao et al., 2023b), the segmentation TTA community also lacks insightful guidelines. For instance, are classic TTA strategies, such as normalization and teacher-student (TS) schemes still effective in segmentation TTA? What is the challenge to address LT problems? Are classic TTA techniques robust to the batch dependency of the test data? What kind of deep architecture is preferred, Transformer or CNN (Zhou et al., 2022)? Moreover, what are the possible solutions to improve segmentation TTA when classic strategies fail to work?
Appendix B Abbreviations
In this paper, the frequently used abbreviations are compiled in Table 9.
Table 9. Abbreviations frequently used in this paper.
Complete description Abbreviation
Teacher-Student TS
Augmentation Aug
Ground-Truth GT
Batch normalization BN
Source only SO
Single-model Single
Pseudo labeling PL
Pseudo-labels PLs
Layer normalization LN
Group normalization GN
Appendix C Transformer-based Architectures are Preferred
In our experiments, we deploy Segformer-B5 (Xie et al., 2021), a Transformer-based architecture, for segmentation TTA tasks. Compared to CNN-based architectures, the backbone of Transformer employs fewer BN layers. We apply DeepLabv3+ (Chen et al., 2018), a typical CNN-based architecture, on datasets ACDC (Sakaridis et al., 2021), Cityscapes-foggy (CS-fog) (Sakaridis et al., 2018) and Cityscapes-rainy (CS-rain) (Hu et al., 2019). The results are depicted in Table 8, where we can observe an obvious drop compared to the results presented in Table 1 of the main manuscript. Thus, it is better to build segmentation TTA architectures based on Transformer instead of CNN. Based on the analysis of normalization updating in Section 3 of the main manuscript, it might be the attention mechanism of Transformer that contributes to its effectiveness in segmentation TTA.
Appendix D More Results Regarding Batch Dependency
Since online adaptation is one of the key characteristics of TTA, we also carry out experiments based on TENT (Wang et al., 2021) besides the single-model and TS scheme. The results are displayed in Table 12, further indicating that TENT is not sensitive to the temporal order of test samples. The reason might be that fewer parameters need to be updated in the deep architecture of TENT, compared to that in the single-model and TS scheme.
Figure 7. Independently calculating the ECE for both correct and incorrect predictions (ACDC-fog, %).
Appendix E More Results under Long-tailed Phenomenon
Although conventional wisdom may suggest that the performance of majority classes surpasses that of minority classes, we observe that this rule does not hold true in segmentation tasks. For example, in the third plot of Figure 8, class 19 attains an IoU of 0.59, whereas class 7 achieves an IoU of 0.52. However, it is worth noting that the count of class 7 is $10^{7}$ while the count of class 19 is $10^{5}$, as illustrated in Figure 12. In summary, a segmentation task in TTA proves to be significantly more intricate than a classification task. The reason might be that the long-tailed (LT) phenomenon may cause error accumulation at the pixel-level and negatively affect the training process. We provide more results on the night, rain, and snow domains within the dataset ACDC, further indicating the complexity of LT problems in segmentation TTA. For instance, after adaptation, the Recall of class 7 increases from 0.27 to 0.68, while the Precision decreases from 0.78 to 0.73. An increase in Recall alongside a decrease in Precision implies a reduction in False Negative and an increase in False Positive. In summary, combining with a region-level solution and introducing data augmentation might be a potential solution to address the LT phenomenon as discussed in Section 5.2 of the main manuscript.
Figure 9. Quantitative metrics analysis on ACDC-night.
Figure 10. Quantitative metrics analysis on ACDC-rain.
Figure 11. Quantitative metrics analysis on ACDC-snow.
Appendix F The Effect of ATTENTION
Our work demonstrates that the attention mechanism plays a pivotal role in a Transformer-based model, which is also shown in Table 3 of the main manuscript. We have shown that GN and LN do not perform well in pixel-level segmentation TTA, as displayed in Figure 1 of the main manuscript. The results demonstrate that updating Normalization layers is not very effective in segmentation TTA while updating the attention mechanism is a promising and novel direction for transformer-based models as illustrated in that Table.
Appendix G State-of-the-art and Recent Segmentation Methods
We use OneFormer (Jain et al., 2023), a typical state-of-the-art and recent segmentation method, as the pre-trained model instead of SegFormer (Xie et al., 2021). As shown in Table 10, although OneFormer shows better performance, it still deteriorates when updating BN layers. We also adopt SAM (Kirillov et al., 2023) and find that it encounters the same problem. These results indicate that our previous analysis is reasonable and solid.
Table 10. Results on two state-of-the-art and recent segmentation approaches, i.e., OneFormer (Jain et al., 2023) and SAM (Kirillov et al., 2023).
Method A-fog A-night A-rain A-snow
OneFormer + SO 70.5 48.7 62.3 61.8
OneFormer + TENT 69.1 46.5 61.2 59.8
OneFormer + SAM + SO 74.9 50.8 64.4 65.1
OneFormer + SAM + TENT 73.8 49.6 64.5 64.1
Table 11. Comparisons between our prompt-based solution (Ours) and other methods related to prompt. It is clear that the performance of Ours is the best.
Method SO DePT DVPT UniPT SVDP Ours
CS (GTA) 68.6 65.1 66.3 60.2 69.1 71.1
CS (Syn) 51.1 48.2 48.6 43.3 52.2 56.1
Appendix H More Results of Model Calibration: Reflecting the Complexity of Segmentation TTA
In the real world, a decision-making system, such as an autonomous car, should not only improve the decision accuracy but also understand when they are potentially unreliable (Guo et al., 2017; Wang et al., 2023a). Attaining an optimal solution in practice proves elusive. Thus, we conduct experiments to delve into model interpretability, aiming to unearth the primary challenges associated with the uncertainty of segmentation TTA where there lacks a comprehensive study on model calibration.
Miscalibration arises from a misalignment between predictive confidence and accuracy, as defined by the expected calibration error (ECE) formalism, i.e., $ECE=\sum_{i=1}^{m}{\frac{|B_{i}|}{N}\left|\text{acc}\left(B_{i}\right)-\text{conf}\left(B_{i}\right)\right|}$, where m is the number of bins, $B_{i}$ denotes a set of samples falling into the bin, and $\textit{acc}\left(B_{i}\right)$ and $\textit{conf}\left(B_{i}\right)$ are actual accuracy and confidence averaged over the samples in the bin, respectively. As displayed in Figure 7, the ECE arising from incorrect predictions markedly outweighs that from correct predictions for both methods. This disparity underscores the predominant role of mispredictions in leading to miscalibration, and it also reinforces the argument that over-confidence remains a paramount concern in segmentation TTA (Wang et al., 2023a).
Table 12. Comparisons under different temporal orders of images from datasets ACDC, Cityscapes-fog, and Cityscapes-rain (%, TENT). Different random seeds (i.e., 0/9/99/999/$9$) represent different time orders. For each row of the table, the results under different random seeds are relatively stable, representing that this approach is not sensitive to the order of test samples.
Domain 0 9 99 999 $9$
A-fog 65.8 65.6 65.6 65.6 65.5
A-night 40.5 41.0 41.1 40.9 41.0
A-rain 62.0 62.2 62.3 62.2 62.0
A-snow 57.8 57.9 57.7 57.9 57.8
CS-fog 73.8 73.8 73.7 73.8 73.8
CS-rain 66.8 66.8 66.8 66.7 66.8
Figure 12. The class distribution in ACDC-fog is highly imbalanced, where the order of magnitude for classes 1 to 9 exceeds $10^{8}$ while that for classes 10 to 19 just exceeds $10^{6}$.
Appendix I Visualization of Segmentation TTA Results
Figure 13. Qualitative comparisons of segmentation results on dataset ACDC. Compared to SO (Source Only), the black box indicates inferior results while the white box signifies improved outcomes.
In this Section, we will visualize the results of different segmentation TTA approaches applied to the dataset ACDC. Some of the results are displayed in Figure 13, where it is clear that TENT (Wang et al., 2021) is hard to differentiate between the road and the sky (marked in black boxes). Moreover, thanks to the TS scheme and the data augmentation strategy, CoTTA (Wang et al., 2022a) produces a more refined segmentation map (shown in white boxes).
Figure 14. Segmentation results of different strategies in CoTTA (Wang et al., 2022a) applied on dataset ACDC. “TS”/“wa”/“sr” are abbreviations for teacher-student scheme/ weight-averaged strategy/stochastic restore, respectively.
Figure 15. Confusion matrix of ACDC-fog. Here, x-axis indicates the predicted labels, while y-axis represents the ground-truth labels. Moreover, the data has been normalized to Min-Max Normalization. We can observe a substantial disparity in performance between the majority and minority classes, underscoring the challenges inherent in segmentation TTA.
The presence of noisy pseudo-labels tends to aggravate error accumulation and catastrophic forgetting in TTA (Wang et al., 2022a; Niu et al., 2022; Yuan et al., 2023). However, we find the experimental results of CoTTA (Wang et al., 2022a) and “SO + aug” are extremely similar, confusing the actual impact of error accumulation and catastrophic forgetting on segmentation TTA. To elucidate this, we conduct a more refined visual analysis, focusing on two strategies proposed by CoTTA (Wang et al., 2022a), i.e., weight-averaged and stochastic restore, As depicted in Figure 14, we can find that these strategies can not guarantee results improvement. For example, in ACDC-fog (shown in the white box), “TS” correctly identifies pixels labeled as sidewalk, although accompanied by numerous misclassifications (the upper part in the box). Utilizing the weight-averaged strategy eliminates these misclassifications, but compromises sidewalk predictions. The subsequent application of the stochastic restore strategy yields prediction in more complex sidewalk areas (the left area in the box) but reintroduces prior noise. A similar pattern is discernible across the remaining domains. In summary, these strategies are not thoroughly effective in genuinely resolving the issues of error accumulation and catastrophic forgetting. Thus, further improvement of segmentation TTA approaches is necessary.